Can I have some help with this one too please? What is the equation of the parabola, with vertex at (2,-4) and directrix y=-6
A.(x+2)^2=8(y+4)
B.(y+4)^2=8(x-2)
C.(y+6)^2=-8(x+2)
D.(x-2)^2=8(y+4)

Found 2 solutions by richwmiller, josgarithmetic:
Answer by richwmiller(17219)

vertex at (2,-4)
(x-h)^2=4p(y-k)
(x-2)^2=4p(y+4)
Without even calculating p the only option that matches is D.(x-2)^2=8(y+4)
and yes p=2

Answer by josgarithmetic(39627)

Vertex given means you want a factor of x-2 and y+4. Only choice D has these.

Otherwise you can find the focus and use definition of parabola and given directrix and found focus and derive the equation using definition of parabola and distance formula.